# Configurations, Tessellations and Tone Networks

### v4

@@ I Introduction
The central argument that we maintain throughout the discussion is that the tonnetz is a mathematical object – in this case, specifically, a geometric configuration – that can be represented by a Levi graph. As an application of our analysis of the cycles of the tonnetz, we take a tentative look in Section IV at the problem of the resolution of the so-called Tristan chord as it arises in excerpts of works by Wagner, Chopin and Tchaikovsky that we exhibit in Figures 5, 6, 7 and 8. Our preliminary analysis of works by these nineteenth century composers involves a reduction from tetradic harmony to triadic harmony. Such a reduction, which can be regarded as a kind of reconnaissance, a survey of the landscape before entering into tetradic territory, can all the same be surprisingly effective and numerous insights can be gained on the basis of this approach.
@@
The idea that tetradic harmonies can be analyzed, tentatively, by use of triadic tonnetz relations can be found in the works of Lewin, Cohn, Tymoczko, and numerous others – and to that extent indeed we but follow here in their footsteps – nonetheless, the idea of using the Levi graph associated with the tonnetz configuration as the fundamental basis for such analysis is entirely new, as best as we can say, and there are three immediate benefits: (a) the discussion can often be simplified and/or made more transparent by use of the Levi graph as a very efficient ‘road map’ of the tonnetz, (b) the geometry of the chord space can be quite unambiguously visualized and surveyed, and (c) the approach can be generalized in a satisfying and unexpected way to tetradic harmonies, which we discuss in Section X.
@@
In Section V we look at pentatonic music. Our starting point is the complete pentagon $\{5_{4},10_{2}\}$ associated with a set of five points in general position in $\mathbb{R}^{2}$, depicted in Figure 9, which leads us in Figures 10 and 11 to show how a tonnetz for pentatonic music can be modelled by use of the self-dual Desargues configuration $\{10_{3}\}$. The associated Levi graph is shown in Figure 12, in two different ways, from which the cycles of the pentatonic system can be deduced. The results are summarized in Proposition 3, which lays the foundation for new types of composition in pentatonic music. A generalization of the idea is presented in Proposition 4, which offers the scope for new compositions in any scale based on an odd number of distinct notes in the octave.
In Section VI we look more closely at the relation between (a) the Eulerian tonnetz, represented by the Levi graph in Figure 2 and (b) the pentatonic tonnetz, represented by the Levi graph in Figure 12. These graphs bear a striking family resemblance, yet the Eulerian tonnetz is usually motivated via the consideration of smooth voice leading, whereas our pentatonic tonnetz was derived via the consideration of set inclusion relations. A critic could argue that if tonnetze are mainly about voice-leading, then the pentatonic construction, interesting as it may be, is not really a tonnetz. We argue the opposite by demonstrating that the Eulerian tonnetz can be derived entirely via set-inclusion relations, as in the pentatonic situation, with no reference to voice leading, minimal work, or parsimony.
@@
Specifically, we show that the incidence relation between the twelve notes of the chromatic scale and the twelve major chords, viewed combinatorially, is that of the D222 of Daublebsky von Sterneck. No mention of the minor chords or their voice-leading relations to major chords are involved in making this assertion. The relation between the notes of the chromatic scale and the minor chords is again that of the D222. The graphs showing the relevant inclusion relations are plotted in Figure 13. These distinct tone-note graphs can be combined with that of the Eulerian tonnetz to produce a grand tripartite tonnetz in the form of a triangular tessellation shown in Figure 14. The result is summarized in Proposition 5.
In Section VII we consider twelve-tone music and show how Sylvester’s combinatorial theory of ‘duads’ and ‘synthemes’ for the permutations of six objects can be used to construct a tonnetz for music based on the specification of an unordered collection of six tones. Indeed, in the context of the system of Schönberg it has long been regarded as natural to consider such hexachords as a basis for both the composition and the analysis of music.
Beginning with a hexachord, one can construct a system comprising fifteen duads (unordered pairs of tones) and fifteen synthemes (unordered triples of non-overlapping duads). These can be represented geometrically by a configuration $\{15_{3}\}$ comprising fifteen lines and fifteen points in $\mathbb{R}^{2}$ described in a paper by Richmond (1900) and discussed in detail by Baker (1925). The Levi graph representing this configuration (Tutte 1947, Coxeter 1957) is shown in Figure 15 and the $\{15_{3}\}$ is drawn in Figure 16. The results are summarized in Proposition 6 .
In Section VIII we return to the analysis of tonal music, with a view to tackling the problem of tetrachords. The strategy, however, is first to look at the relation between major and minor triads when exactly two tones are allowed to change in a chord transition. We show that the resulting bipartite graph has two components, each containing six major chords and six minor chords.
These are shown in Figure 17. Each component has girth four: hence neither is a Levi graph – but each component can be unfolded to yield the infinite graph of an Archimedean or semi-regular tessellation – a tiling of the plane by hexagons, squares, and dodecagons, for which one component is shown in Figure 18. The resulting cycles include some interesting sonorities and in Figure 19 we have composed a short fanfare based on one of the hexacycles of this tonnetz.
@@
In Section IX we take a break from the main argument and present a short dialogue involving three friends, Salviati, Sagredo and Simplicio, who are interested in mathematics and music. They have agreed to meet for drinks at a music cafe in central London, but Simplicio is running late. Salviati and Sagredo enter into a conversation about the tonnetz and its mathematical representations. Then in Section X we look at the Tristan genus, comprising the twelve dominant sevenths and the twelve minor sixths. Previous work in this area (Childs 1998, Douthart & Steinbach 1998, Tymoczko 2011, Cohn 2012) has tended to look at irregular graphs, with the dominant sevenths and the half-diminished sevenths as vertices, and the edge sets including all possible ‘parsimonious’ transformations, or some large subset thereof, variously supplemented by further vertices representing additional, less consonant tetrachords, such as the six French sixths and the three diminished sevenths, and corresponding edges. The irregularities of these graphs, the large numbers of edges, the heterogeneity of the vertex constituents make one wonder whether one can do better.
@@
We consider the situation when exactly two tones are allowed to change in a chord transition between two chords of the reductively opposite mode when we regard the minor sixth chord as a minor triad with an added dissonant sixth tone and regard the major seventh chord as a consonant major triad with an added flat seventh tone. This gives a six to six map between the two sets of twelve chords. But this cannot give rise to a configuration, since the fundamental inequality (1) is violated. The inequality tells us that to obtain a bipartite graph of degree six with no cycles of degree less than six, the number of vertices must be no less than thirty-one – but we only have twenty four. To obtain a tractable subclass of transformations that leads to a geometry, we must cut down the number of allowed transformations. Our proposed solution to the problem, which is new as far as we can see, is as follows: we constrain the transformation class in such a way that it always retains the dissonant note and hence always changes two tones of the underlying major or minor chord. Note that the image of the dissonant tone need not (but might) be a dissonant tone of the transformed chord. The result is a three to three map between the dominant sevenths and the minor sixths with no tetracycles – a configuration $\{12_{3}\}$. We draw its Levi graph in Figure 20. One can tell at a glance that this Levi graph is distinct from that of the Eulerian tonnetz by counting the numbers of cycles of various lengths. A single discrepancy suffices to show that our Tristan tonnetz is different from the Eulerian one. But all the configurations $\{12_{3}\}$ are known, so if it is not the Eulerian tonnetz, which one is it? Surprisingly, it turns out that the Tristan tonnetz is the D228 of Daublebsky von Sterneck, which we show – with tetrachord labels attached – in Figure 21. This configuration is as memorable as the D222 and it has a large group of symmetries. We conclude Section X with an analysis of passages in Tristan, Götterdämmerung and Parsifal that allow for comparison of our method with arguments of Lewin (1996), Tymoczko (2011) and Cohn (2012). For the Finale in Section XI we return to the Fidelio Cafe, where Simplicio as at last arrived.
@@ II The Tonnetz as a Levi Graph
It is probably fair to say that most investigations of the tonnetz hitherto have tended to develop the chicken-wire representation put forward by Douthett & Steinbach (1998) and the group-theoretic ideas that are implicit in this representation. Though not inconsistent with that point of view, we prefer a somewhat different approach to the tonnetz that allows more readily for the far-reaching generalizations required for treating forms of music other than that of the extended common practice period. Our starting point is the observation that the tonnetz can be represented by a bipartite graph of the form shown in Figure 2. Major chords are represented by white vertices and minor chords are represented by black vertices. Each major chord is connected to the vertices of three minor chords and each minor chord is connected to the vertices of three major chords. For example, $CM$ is connected to $Em$, CmCm, AmAm, whereas $Cm$ is connected to $E_{\flat}M$, $CM$, $A_{\flat}M$.
Various well-known features of the tonnetz can be read off from this graph and less well-known aspects of the tonnetz also become apparent. To develop our approach further, we recall a few ideas from graph theory (Harary 1969, Wilson 1972, Bondy & Murty 1976). We write $\mathbb{N}=\{1,2,3,.\,.\,.\,\}$ for the natural numbers and $\mathbb{Z}^{+}=\{0,1,2,.\,.\,.\,\}$ for the non-negative integers  [^8] There are diverse notations for these basic sets. Here we follow Welsh 1976 and Williams 1991.. A graph is a triple $G=(V,E,\pi)$ where $V$ is a non-empty finite set of elements called vertices, $E$ is a set of elements called edges and $\pi:E\to\mathcal{P}(V)$ is a map from $E$ to the set $\mathcal{P}(V)=\{\{u,v\}:u,v\in V\}$. We observe that $\mathcal{P}(V)$ is the set of unordered pairs of elements of $V$. Care is required when we consider elements of $\mathcal{P}(V)$ of the form $\{v,v\}$, $v\in V$. Since $\{v,v\}=\{v\}$ as sets, we have $\mathcal{P}(V)=\{\{u,v\}:u,v\in V,u\neq v\}\cup\{\{v\}:v\in V\}$. If $\pi(e)=\{v_{1},v_{2}\}$, where $v_{1},v_{2}\in V$ and $e\in E$, we say that $v_{1}$ and $v_{2}$ are joined by $e$ and that $e$ is incident with $v_{1}$ and $v_{2}$.
@@
The number of edges joining a vertex is the degree of that vertex. The degree of a graph is the maximum of the degrees of its vertices. If all the vertices have the same degree, the graph is regular. If the edges $e_{1}$ and $e_{2}$ are incident with a common vertex, we say that $e_{1}$ and $e_{2}$ are adjacent. If a vertex is incident with no edges it is isolated. If $G$ has no loops or multiple edges it is simple. A sequence of $k\in\mathbb{N}$ distinct consecutively adjacent edges is a path of length $k$. A graph is connected if any two vertices are joined by some path. A subgraph of $G=(V,E,\pi)$ is a graph $G^{\prime}=(V^{\prime},E^{\prime},\pi^{\prime})$ such that $V^{\prime}\subset V$, $E^{\prime}\subset E$, $E^{\prime}=\pi^{-1}(\mathcal{P}(V^{\prime}))$, and $\pi^{\prime}$ is the restriction of $\pi$ to $E^{\prime}$. If $G$ is not connected then it splits into two or more connected subgraphs called components of $G$.
A closed sequence of $k$ distinct consecutively adjacent edges that begins at some vertex and comes back to that vertex is called a $k$-cycle. A loop is a $1$-cycle; a pair of multiple edges form a $2$-cycle. The least value of $k$ for which a graph $G$ admits a $k$-cycle is called the girth of $G$. If $V$ has $N$ elements then $N$ is the order of the graph. If the vertices of a regular graph are of degree $d$, then $dN=2M$, where $M$ is the number of edges. By a forest we mean a graph with no cycles. By a tree we mean a connected forest. Finally, we say that a graph is bipartite if every cycle has an even number of edges, in which case there exist sets $V_{0}(G)$ and $V_{1}(G)$ such that $V(G)=V_{0}(G)\cup V_{1}(G)$ with $V_{0}(G)\cap V_{1}(G)=\varnothing$.
@@
Figure 2: The tonnetz as a Levi graph. The tonnetz can be modelled as a regular bipartite graph of degree three with twelve white vertices corresponding to major chords and twelve black vertices corresponding to minor chords. Each major chord is linked to three minor chords and each minor chord is linked to three major chords. The four $3p$-hexacycles appear as blunted triangles such as $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$. There are twelve $2p$-hexacycles (bow ties), one beginning at each major chord, a typical example being $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$. Each major chord is contained in three $2p$-hexacycles; similarly for minor chords.
By a combinatorial configuration of type $\{m_{r},n_{k}\}$ for $m,n,r,k\in\mathbb{N}$ we mean an incidence structure with $m$ points and $n$ lines such that $k$ points lie on each line and $r$ lines go through each point. We use the term ‘type’ here since in some cases several different configurations of the same type may exist. By a flag we mean a pair $\{p,L\}$ where $p\in\mathbb{P}$ and $L\in\mathbb{L}$. One then easily sees that for any configuration $\{m_{r},n_{k}\}$ we have the fundamental equality $mr=nk$, since each side of this equation gives a way of counting the total number of distinct flags in the configuration. The fundamental inequality
@@
The significance of Levi graphs is that their correspondence with combinatorial configurations is one to one (Grünbaum 2009, p. 28). To wit: a graph $G$ is the Levi graph of an $\{m_{r},n_{k}\}$ combinatorial configuration if and only (i) $G$ is bipartite with partition $V(G)=P(G)\cup L(G)$, where $P$ has cardinality $m$ and $L$ has cardinality $n$; (ii) every vertex in $P$ is of degree $r$ and every vertex in $L$ is of degree $k$; (iii) $G$ is simple; (iv) $G$ has girth at least six. With these definitions and facts at hand, we are able to assert the following:
@@
In particular, the tonnetz is a combinatorial configuration. Inspection of Figure 2 shows the minimal cycles are hexacycles. Hexacycles can be classified by the number of parallel major-minor transformations that they contain, that is to say, transformations of the form $CM$ to $Cm$ or $Cm$ to $CM$. Figure 2 shows that there are exactly four hexacycles admitting three such parallel transformations, each taking the form of a sort of blunted triangle. These four distinctive hexacycles partition the tonnetz and are given as follows:
$\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle,\hskip 14.22636pt\langle FM,Fm,D_{\flat}M,C_{\sharp}m,AM,Am,FM\rangle,\hskip 14.22636pt$
@@
Figure 3: We illustrate a $3p$-hexacycle, a $2p$-hexacycle, a $4p$-octacycle and a $2p$-decacycle. The $3p$-hexacycle $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$ is one of the four basic hexacycles discussed by Cohn and numerous other authors. There are twelve overlapping ‘straight bow tie’ $2p$-hexacycles. We see $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$ as an example. Similarly, we have the $4p$-octatonic ‘four-cornered hat’ cycle $\langle CM,Am,AM,F_{\sharp}m,G_{\flat}M,E_{\flat}m,E_{\flat}M,Cm,CM\rangle$ and the $2p$-decatonic ‘floppy bow tie’ cycle $\langle CM,Am,FM,Dm,B_{\flat}M,B_{\flat}m,D_{\flat}M,Fm,A_{\flat}M,Cm,CM\rangle$.
The edges of the tonnetz fall into three sets of twelve. In the first set we have edges of the form $\{CM,cm\}$, which we already mentioned, that transform a major (resp., minor) chord into its parallel minor (resp., major). In the second set we have edges of the form $\{CM,Am\}$ that transform a major (resp., minor) chord into its relative minor (resp., major). In the third set we have edges of the form $\{CM,Em\}$ that transform a major (resp., minor) chord into a corresponding minor (resp., major) under a so-called leading tone exchange. Three edges meet at each vertex, one of each type.
One can ask if there are any other hexacycles in the Levi graph of the tonnetz apart from the four hexatonic hexacycles and interestingly the answer is affirmative. There are twelve further hexacycles, one beginning at each major chord, of which a typical example is

### v5

@@ I Introduction
A central argument that we maintain throughout the discussion is that the tonnetz is a mathematical object – in this case, specifically, a combinatorial configuration – that can be represented by a Levi graph. Thus, a certain abstract structure manifests itself as (i) a combinatorial configuration, (ii) a geometric configuration – and (iii) a ‘musical’ configuration, the latter being of course the tonnetz. As an application of our analysis of the cycles of the tonnetz, we take a look in Section IV at the problem of the resolution of the ‘Tristan’ chord as it arises variously in works by Wagner, Chopin and Tchaikovsky, excerpts of which we exhibit in Figures 5, 6, 7 and 8. Our preliminary analysis of compositions by these nineteenth century composers involves a reduction from tetradic harmony to triadic harmony. Such a reduction, which can be regarded as a kind of reconnaissance, a survey of the landscape before entering into tetradic territory, can all the same be surprisingly effective and numerous insights can be gained on the basis of this approach.
@@
The idea that tetradic harmonies can be analyzed, albeit tentatively, by use of triadic tonnetz relations can be found in the works of Lewin, Cohn, Tymoczko, and numerous other authors – and to that extent indeed we but follow in their footsteps – nonetheless, the idea of using the Levi graph associated with the tonnetz configuration as the fundamental basis for such analysis is entirely new, as best as we can say, and there are three immediate benefits: (a) the discussion can often be simplified and/or made more transparent by use of the Levi graph as a very efficient ‘road map’ of the tonnetz, (b) the geometry of the chord space can be quite unambiguously visualized and surveyed, and (c) the approach can be generalized in a satisfying and unexpected way to tetradic harmonies, which we discuss in Section X.
@@
In Section V we consider aspects of pentatonic music. Our starting point is the complete pentagon $\{5_{4},10_{2}\}$ associated with a set of five points in general position in $\mathbb{R}^{2}$, depicted in Figure 9, which leads us in Figures 10 and 11 to show how a tonnetz for pentatonic music can be modelled by use of the self-dual Desargues configuration $\{10_{3}\}$. The associated Levi graph is shown in Figure 12, in two different ways, from which the cycles of the pentatonic system can be deduced. The results are summarized in Proposition 3, which lays a foundation for new types of composition in pentatonic music. A generalization of the idea is presented in Proposition 4, which offers the scope for new compositions in any scale based on an odd number of distinct notes in the octave.
@@
In Section VI we look more closely at the relation between the Eulerian tonnetz, represented by the Levi graph in Figure 2, and the pentatonic tonnetz, represented by the Levi graph in Figure 12. These graphs bear a striking resemblance, yet whereas construction of the Eulerian tonnetz is usually motivated by the consideration of smooth voice leading, our pentatonic tonnetz was derived via set inclusion relations. One might try to argue that if tonnetze are mainly about voice-leading, then the pentatonic construction is not really a tonnetz. We argue the opposite by demonstrating that the Eulerian tonnetz itself can be derived entirely via the consideration of set-inclusion relations, as in the pentatonic situation, with no reference to voice leading, minimal work, or parsimony. Specifically, we show that the incidence relation between the twelve notes of the chromatic scale and the twelve major chords, viewed combinatorially, is that of the D222 of Daublebsky von Sterneck. No mention of the minor chords or their voice-leading relations to major chords is involved in making this assertion. The relation between the notes of the chromatic scale and the minor chords is again that of the D222. The graphs showing the relevant inclusion relations are plotted in Figure 14. These distinct note/chord graphs can be combined with that of the Eulerian tonnetz to produce a grand tripartite tonnetz given by the tessellation shown in Figure 15. The results are summarized in Proposition 5.
In Section VII we consider twelve-tone music and show how Sylvester’s combinatorial theory of ‘duads’ and ‘synthemes’ for the permutations of six objects can be used to construct a tonnetz for music based on the specification of an unordered collection of six tones. Indeed, in the context of the system of Schönberg it has long been regarded as natural to consider such hexachords as a basis for both the composition and the analysis of music. Beginning with a hexachord, one can construct a system consisting of fifteen duads (unordered pairs of tones) and fifteen synthemes (unordered triples of non-overlapping duads). These can be represented geometrically by a configuration $\{15_{3}\}$ comprising fifteen lines and fifteen points in $\mathbb{R}^{2}$ described in a paper by Richmond (1900) and discussed in detail by Baker (1925). The Levi graph representing this configuration (Tutte 1947, Coxeter 1957) is shown in Figure 16 and a geometric representation of this $\{15_{3}\}$ is drawn in Figure 17. The results are summarized in Proposition 6. In Section VIII we return to the analysis of tonal music, with a view to tackling the problem of tetrachords. The strategy, however, is first to look at the relation between major and minor triads when exactly two tones are allowed to change in a chord transition. We show that the resulting bipartite graph has two components, each containing six major chords and six minor chords. These are shown in Figure 19. Each component has girth four: hence neither is a Levi graph – but each component can be unfolded to yield the infinite graph of an Archimedean or semi-regular tessellation – a tiling of the plane by hexagons, squares, and dodecagons, for which one component is shown in Figure 20. The resulting cycles include some interesting sonorities and in Figure 21 we present a short fanfare that we have composed based on one of the hexacycles of this tonnetz.
@@
In Section IX we take a break from the main argument and present a short dialogue involving three friends, Salviati, Sagredo and Simplicio, who are interested in mathematics and music. They have agreed to meet for drinks at a music cafe in central London, but Simplicio is running late. Salviati and Sagredo enter into a conversation about the tonnetz and its mathematical representations. Then in Section X we look at the Tristan genus  [^9] We use the terminology of Cohn (2012), pp. 159-166. comprising the twelve dominant sevenths and the twelve minor sixths. Previous work in this area (Childs 1998, Douthart & Steinbach 1998, Tymoczko 2011b, Cohn 2012) has tended to look at irregular graphs, with the dominant sevenths and the half-diminished sevenths as vertices, and the edge sets including all possible ‘parsimonious’ transformations, or some large subset thereof, variously supplemented by further vertices representing additional, less consonant tetrachords, such as the six French sixths and the three diminished sevenths, and corresponding edges. The irregularities of these graphs, the large numbers of edges, the heterogeneity of the vertex constituents make one wonder whether one can do better.
@@
We consider the situation when exactly two tones are allowed to change in a chord transition between two chords of the reductively opposite mode when we regard the minor sixth chord as a minor triad with an added dissonant sixth tone and we regard the major seventh chord as a consonant major triad with an added flat seventh tone. This gives a six to six map between the two sets of twelve chords. But this cannot give rise to a configuration, since the fundamental inequality (1) is violated. The inequality tells us that to obtain a bipartite graph of degree six with no cycles of degree less than six, the number of vertices must be no less than thirty-one – but we only have twenty four. To obtain a tractable subclass of transformations that leads to a geometry, we must cut down the number of allowed transformations. Our proposed solution to the problem, which is new as far as we can see, is as follows: we constrain the transformation class in such a way that it always retains the dissonant note and hence always changes two tones of the underlying major or minor chord. Note that the image of the dissonant tone need not (but might) be a dissonant tone of the transformed chord. The result is a three to three map between the dominant sevenths and the minor sixths with no tetracycles – a configuration $\{12_{3}\}$. We draw its Levi graph in Figure 22. One can tell at a glance that this Levi graph is distinct from that of the Eulerian tonnetz by counting the numbers of cycles of various lengths. A single discrepancy suffices to show that our Tristan tonnetz is different from the Eulerian one. But all the configurations $\{12_{3}\}$ are known, so if it is not the Eulerian tonnetz, which one is it? Surprisingly, it turns out that the Tristan tonnetz is the D228 of Daublebsky von Sterneck, which we show – with tetrachord labels attached – in Figure 23. This configuration is as memorable as the D222 and it has a large group of symmetries. We conclude Section X with an analysis of passages in Tristan, Götterdämmerung and Parsifal that allow for comparison of our method with arguments of Lewin (1996), Tymoczko (2011b) and Cohn (2012). For the Finale in Section XI we return to the Fidelio Cafe, where Simplicio has at last arrived.
@@ II The Tonnetz as a Levi Graph
Most investigations of the tonnetz hitherto have tended to develop the chicken-wire representation put forward by Douthett & Steinbach (1998) and the group-theoretic ideas that are implicit in this representation. Though not inconsistent with that point of view, we prefer a different approach to the tonnetz that allows more readily for the far-reaching generalizations required for treating forms of music other than that of the extended common practice period.
Our starting point is the observation that the tonnetz can be represented by a regular bipartite graph of degree three and order 24 of the form shown in Figure 2. Major chords are represented by white vertices and minor chords are represented by black vertices. Each major chord is connected to the vertices of three minor chords and each minor chord is connected to the vertices of three major chords. For example, $CM$ is connected to $Em$, CmCm, AmAm, whereas $Cm$ is connected to $E_{\flat}M$, $CM$, $A_{\flat}M$.
Various well-known features of the tonnetz can be read off from this graph and less well-known aspects of the tonnetz also become apparent. To develop our approach further, we recall a few ideas from graph theory (Harary 1969, Wilson 1972, Bondy & Murty 1976). We write $\mathbb{N}=\{1,2,3,.\,.\,.\,\}$ for the natural numbers and $\mathbb{Z}^{+}=\{0,1,2,.\,.\,.\,\}$ for the non-negative integers  [^10] There are diverse notations for these basic sets. We follow Welsh (1976) and Williams (1991).. A graph is a triple $G=(V,E,\pi)$ where $V$ is a non-empty finite set of elements called vertices, $E$ is a set of elements called edges and $\pi:E\to\mathcal{P}(V)$ is a map from $E$ to the set $\mathcal{P}(V)=\{\{u,v\}:u,v\in V\}$. We observe that $\mathcal{P}(V)$ is the set of unordered pairs of elements of $V$. Care is required when we consider elements of $\mathcal{P}(V)$ of the form $\{v,v\}$, $v\in V$. Since $\{v,v\}=\{v\}$ as sets, we have $\mathcal{P}(V)=\{\{u,v\}:u,v\in V,u\neq v\}\cup\{\{v\}:v\in V\}$. If $\pi(e)=\{v_{1},v_{2}\}$, where $v_{1},v_{2}\in V$ and $e\in E$, we say that $v_{1}$ and $v_{2}$ are joined by $e$ and that $e$ is incident with $v_{1}$ and $v_{2}$.
@@
The number of edges joining a vertex is the degree of that vertex. The degree of a graph is the maximum of the degrees of its vertices. If all the vertices have the same degree, the graph is regular. If the edges $e_{1}$ and $e_{2}$ are incident with a common vertex, we say that $e_{1}$ and $e_{2}$ are adjacent. If a vertex is incident with no edges it is isolated. If $G$ has no loops or multiple edges it is said to be simple. A sequence of $k\in\mathbb{N}$ distinct consecutively adjacent edges is a path of length $k$. A graph is connected if any two vertices are joined by some path. A subgraph of $G=(V,E,\pi)$ is a graph $G^{\prime}=(V^{\prime},E^{\prime},\pi^{\prime})$ such that $V^{\prime}\subset V$, $E^{\prime}\subset E$, $E^{\prime}=\pi^{-1}(\mathcal{P}(V^{\prime}))$, and $\pi^{\prime}$ is the restriction of $\pi$ to $E^{\prime}$. If $G$ is not connected then it splits into two or more connected subgraphs called components of $G$.
A closed sequence of $k$ distinct consecutively adjacent edges that begins at some vertex $v$ and comes back to $v$ in such a way that all $k$ of the vertices are distinct is called a $k$-cycle. A loop is a $1$-cycle; a pair of multiple edges form a $2$-cycle. The least value of $k$ for which a graph $G$ admits a $k$-cycle is called the girth of $G$. If $V$ has $N$ elements then $N$ is the order of the graph. If the vertices of a regular graph are of degree $d$, then $dN=2M$, where $M$ is the number of edges. By a forest we mean a graph with no cycles. By a tree we mean a connected forest. Finally, we say that a graph is bipartite if there exist sets $V_{0}(G)$ and $V_{1}(G)$ satisfying $V(G)=V_{0}(G)\cup V_{1}(G)$ and $V_{0}(G)\cap V_{1}(G)=\varnothing$ such that every edge of $G$ joins a vertex of $V_{0}$ to a vertex of $V_{1}$. One can show that a graph is bipartite if and only if every cycle has an even number of edges.
@@
Figure 2: The Eulerian tonnetz as a Levi graph. The tonnetz can be modelled as a regular bipartite graph of degree three with twelve white vertices corresponding to major chords and twelve black vertices corresponding to minor chords. Each major chord is linked to three minor chords and each minor chord is linked to three major chords. The four $3p$-hexacycles are blunted triangles such as $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$. There are twelve $2p$-hexacycles (bow ties), one beginning at each major chord, an example being $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$. Each major chord is contained in three $2p$-hexacycles; similarly for minor chords.
By a combinatorial configuration of type $\{m_{r},n_{k}\}$ for $m,n,r,k\in\mathbb{N}$ we mean an incidence structure with $m$ points and $n$ lines such that $k$ points lie on each line and $r$ lines go through each point. We use the term ‘type’ here since in some cases several different configurations of the same type may exist. By a flag we mean a pair $\{p,L\}$ where $p\in\mathbb{P}$ and $L\in\mathbb{L}$.
For any configuration $\{m_{r},n_{k}\}$ we have the fundamental equality $mr=nk$, since each side of this equation gives a way of counting the total number of distinct flags in the configuration. The fundamental inequality
@@
The significance of Levi graphs is that their correspondence with combinatorial configurations is one to one (Grünbaum 2009, p. 28). To wit: a graph $G$ is the Levi graph of a combinatorial configuration $\{m_{r},n_{k}\}$ if and only (i) $G$ is bipartite with partition $V(G)=P(G)\cup L(G)$, where $P$ has cardinality $m$ and $L$ has cardinality $n$; (ii) every vertex in $P$ is of degree $r$ and every vertex in $L$ is of degree $k$; (iii) $G$ is simple; (iv) $G$ has girth at least six. With these definitions and facts at hand, we are able to assert the following:
@@
Inspection of Figure 2 shows the minimal cycles are hexacycles. Hexacycles can be classified by the number of parallel transformations that they contain, that is, transformations of the form $CM$ to $Cm$ or $Cm$ to $CM$. Figure 2 shows that there are exactly four hexacycles admitting three such parallel transformations, each taking the form of a sort of blunted triangle. These four distinctive hexacycles partition the tonnetz and are given as follows:
$\hskip 8.5359pt\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle,\hskip 14.22636pt\langle FM,Fm,D_{\flat}M,C_{\sharp}m,AM,Am,FM\rangle,\hskip 14.22636pt$
@@
Figure 3: We illustrate a $3p$-hexacycle, a $2p$-hexacycle, a $4p$-octacycle and a $2p$-decacycle. The $3p$-hexacycle $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$ is one of the four basic hexatonic sequences discussed by Cohn and numerous other authors. There are twelve overlapping ‘straight bow tie’ $2p$-hexacycles. We see $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$ as an example. Similarly, we have the ‘four-cornered hat’ $4p$-octacycle $\langle CM,Am,AM,F_{\sharp}m,G_{\flat}M,E_{\flat}m,E_{\flat}M,Cm,CM\rangle$ and the ‘floppy bow tie’ $2p$-decacycle $\langle CM,Am,FM,Dm,B_{\flat}M,B_{\flat}m,D_{\flat}M,Fm,A_{\flat}M,Cm,CM\rangle$.
The edges of the tonnetz fall into three sets of twelve. In the first set we have edges of the form $\{CM,Cm\}$, which we already mentioned, that transform a major (resp., minor) chord into its parallel minor (resp., major). In the second set we have edges of the form $\{CM,Am\}$ that transform a major (resp., minor) chord into its relative minor (resp., major). In the third set we have edges of the form $\{CM,Em\}$ that transform a major (resp., minor) chord into a corresponding minor (resp., major) under a leading tone exchange. Three edges meet at each vertex, one of each type.
One can ask if there are any other hexacycles in the Levi graph of the tonnetz apart from the four hexatonic hexacycles and the answer is affirmative. There are twelve further hexacycles, one beginning at each major chord, of which a typical example is